OpenStudy (anonymous):

can you solve this for me? i did the rest of the question. 3/10 ---- 4/5

OpenStudy (anonymous):

Do you know what happens when you divide by a fraction?

OpenStudy (anonymous):

you switch the numbers, so 6/5 would become 5/6

OpenStudy (owlcoffee):

We cannot solve it for you but rather guide you through the question. Since it is labelled "high" that means you have no problem in solving the fraction in a more theoretical manner. The division of fractions can be operated if you flip over one of them and turn it into a product: \[\frac{ \frac{ a }{ b } }{ \frac{ x }{ y } } \iff (\frac{ a }{ b })(\frac{ y }{ x })\]

OpenStudy (anonymous):

that means in order to solve it you would do 3/10 x 5/4. this is 3/8 after it has benn reduced, right?

OpenStudy (owlcoffee):

That's correct.

OpenStudy (anonymous):

ohhhh, i found my mistake, thank you!

OpenStudy (owlcoffee):

no problem.